OpenStudy (anonymous):

Find P(Y >1|Y<=2). Y is a continuous random variable. F(y) = {0 when y>0} and {1-e^(-y^2) when y>=0}. f(y) = {0 when y>0} and {2y*e^(-y^2) when y>=0} I know that P(Y>1|Y<=2) = [P(Y<1 AND Y <=2)]/P(Y<=2)] I am confused at how to integrate the >1 part. is that from 1 to 2 or since it is not >= can we not just integrate from 1 to 2?

OpenStudy (anonymous):

on the second paragraph that should be a P(Y>1 not P(Y<1

OpenStudy (anonymous):

also, I was wondering is P(Y>=.5 AND Y>=.3) the same as P(Y>=.5)

OpenStudy (anonymous):

If you guys have some time I would appreciate any help. thanks in advance. @KingGeorge @satellite73

OpenStudy (anonymous):

i am confused by the inequalities, but may be it is because it is late

OpenStudy (anonymous):

but also i am confused why you have to integrate, you have \(F\) maybe i am just confused and should shut up, but can't you compute \[F(2)-F(1)\]?

OpenStudy (anonymous):

do you know if P(Y>1 AND Y <=2) = P(1<Y<=2)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but i don't understand the \(y>0\) and \(y\geq 0\) part

OpenStudy (anonymous):

ok, then since its continuous, I think we can say P(1<Y<=2) = P(1<Y<2)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

the y>0 and y>=0 is different conditions for the F(y) function. I think it is the "parameters" for the range of the function.

OpenStudy (anonymous):

it is \[F(y) = \left\{\begin{array}{rcc} 0 & \text{if} & y<0 \\ 1-e^{-y^2}& \text{if} & y\geq 0 \end{array} \right.\]

OpenStudy (anonymous):

ok then let me shut up because i am confused by that

OpenStudy (anonymous):

yes. I'm pretty bad at writing it in nice formatting.

OpenStudy (anonymous):

i don't understand how it can be one condition for \(y>0\) and another for \(y\geq 0\)

OpenStudy (anonymous):

that didn't make sense to me either. I think my prof might have made a mistype for the 0<y and it should have been 0>y.

OpenStudy (anonymous):

i certainly hope so because it doesn't make sense in any context to me, probability or otherwise

OpenStudy (anonymous):

F(y) is the probability distribution function, while f(y) is the probability density function. for the continuous random variable Y

OpenStudy (anonymous):

i would not bet money on my answer, but my guess is it is \[\frac{F(2)-F(1)}{F(2)-F(0)}\]

OpenStudy (anonymous):

for the numerator of the conditional probability, I am integrating form 1 to 2 over the function \[2y*e ^{-y^2}\]

OpenStudy (anonymous):

*from

OpenStudy (anonymous):

i thought you were given \(F\)

OpenStudy (anonymous):

no matter you will get the same thing

OpenStudy (anonymous):

since the anti derivative is \(-e^{-y^2}\)

OpenStudy (anonymous):

i am getting .35/.35 which equals 1. Now that i think about it it makes a lot of sense. we are taking the probability of everything above 1 and the probability of everything less than or equal to 2 so that would be the probability over all possible values which would be 1.

OpenStudy (anonymous):

less = *else

OpenStudy (anonymous):

i get \[\frac{e^{-1}-e^{-4}}{1-e^{-4}}\]

OpenStudy (anonymous):

probably more elegantly written as \[\frac{e^3-1}{e^4-1}\]

OpenStudy (anonymous):

yeah, that's what i get too. let me check my algebra again.

OpenStudy (anonymous):

yeah i messed up on my algebra for the denominator. instead of .35 it should have been .98.

OpenStudy (anonymous):

or if you want to be real fancy write \[\frac{1+e+e^2}{1+e+e^2+e^3}\]

OpenStudy (anonymous):

since it's a probability, I have to write it as a decimal. I get .357 when i divide.

OpenStudy (anonymous):

such a nice answer such an ugly decimal

OpenStudy (anonymous):

that's true. thank's for all the help. I really appreciate it. now off to sleep...